OpenStudy (anonymous):

The formula for determining interest compounded monthly is A = P(1 + )12t, where A represents the amount invested after t years, P the principal invested, and r the interest rate. Jimmy invests $1,000 at an interest rate of 10% for 3 years, while Jenny invests $1,000 at an interest rate of 5% for 6 years. Determine the amount of return gained by Jimmy and Jenny. In complete sentences, summarize your results.

OpenStudy (anonymous):

formula is \[P(1+\frac{r}{12})^{12t}\] of you are compounding monthly

OpenStudy (anonymous):

10% for three years would therefore be \[1000(1+\frac{.1}{12})^{36}\] and 5% for 10 years would be \[1000(1+\frac{.05}{12})^{120}\] now a calculator to compare

OpenStudy (anonymous):

?......

OpenStudy (kropot72):

The method detailed by satellite73 is correct. First you need to calculate the term inside the brackets: 1 + 0.1/12 = 1.008333. Do you follow this step?